OpenStudy (anonymous):

find the sum in simplest form. 106 / v^6+v - 28/v^6+v

OpenStudy (anonymous):

\[\frac{ 106 }{ v^6+v }-\frac{ 28 }{ v^6+v }\]

OpenStudy (anonymous):

is it \[\frac{ 78 }{ v^6+v }\]

Directrix (directrix):

Yes. I wonder if you are supposed to leave the denominator in factored form. That would give 78 up top and v(v^5 + 1) downstairs Do you have answer options?

Directrix (directrix):

@Alice2112

Directrix (directrix):

The site crashed. Did you finish this?